an 11% salt solution was mixed with an 18% salt solution to make 20L of a 16% solution. How many liters of 11% salt solution were required?
Answer by jorel1380(3719)

m+n=20
.11m+.18n=20(.16)=3.2
11m+18n=320
11m+11n=220
7n=100
n=100/7
m=40/7
You need 40/7 liters of 11% salt solution for your mix..